# LIA: Supervised Fine-Tuning of Large Language Models for Automatic Issue Assignment

### v1

@@ 5. Discussion
Another key finding is that LIA consistently outperformed state-of-the-art graph-based approaches, including NCGBT (dong2024neighborhood). This is notable because graph-based methods have traditionally been strong baselines for issue assignment due to their ability to leverage project structure and developer–issue relationships. In addition, LIA also significantly outperformed our text-based baseline (CBR), showing that a fine-tuned LLM can capture deeper relational patterns between developers and issue semantics, even without relying on explicit graph structures.
@@ 6. Threats to Validity
We provide the entire experiment, including the source code for LIA and scripts evaluation online (ReplicationPackage).
@@ 7. Related Work
Automatic issue assignment has become an important component in modern software maintenance, aiming to reduce manual effort and speed up the resolution of reported defects (dipongkor2023comparative; lee2022light). Over time, researchers have proposed a variety of techniques to automate this process. In this section, we group previous work into two categories: text classification-based and graph-based methods.
@@ 7.1. Text-Based Methods
Text-based approaches have commonly been used for issue assignment by treating parts of issue reports, such as titles and descriptions, as structured input. These inputs are turned into feature representations, and various models are trained to predict the appropriate developer as a classification label. Early work in this space used traditional ML techniques such as Naive Bayes (murphy2004automatic) and Support Vector Machines (SVM) (anvik2006automating), coupled with methods like TF-IDF (xuan2014towards) and Latent Dirichlet Allocation (LDA) (xia2016improving). These models, however, lacked the semantic understanding needed to capture the complex context often found in issue report text (cao2025complex; zhou2025issuecourier; zhang2023ealink; arnob2025bug).
@@
With the development of DL, various classification techniques have been introduced to enhance issue assignment. Wang et al. (wang2024empirical) compared 35 DL-based issue assignment models, combining different text embeddings with various classifiers. They reported that Bi-LSTM models using attention or ELMo embeddings consistently outperformed alternatives in both top-k accuracy and mean reciprocal rank (MRR). Similarly, Mani et al. (mani2019deeptriage) developed a model combining a bidirectional recurrent neural network with attention, using Word2Vec (mikolov2013efficient) to represent text and learn both syntactic and semantic features. Aung et al. (aung2022multi) introduced a multi-task learning model that simultaneously assigned developers and classified issue types. Their model used a CNN for encoding issue text and a BiLSTM for abstract syntax trees (ASTs), allowing it to incorporate structural code features often ignored by previous approaches. Arnob et al. (arnob2025bug) used XLNet for issue assignment and demonstrated that combining commit messages with summaries and descriptions led to better performance by providing richer context than text alone.
@@
More recent studies have applied transformer models that are better at capturing contextual and semantic patterns than the mentioned DL approaches. Lee et al. (lee2022light) proposed a lightweight framework that fine-tunes pretrained transformers like RoBERTa (liu2019roberta) and DeBERTa (he2020deberta) on issue report data, achieving competitive performance with lower computational overhead. Dipongkor et al. (dipongkor2023comparative) evaluated transformer-based models across multiple open-source datasets and found that DeBERTa outperformed BERT (devlin2019bert) and RoBERTa due to its deeper attention mechanisms and disentangled representations. Their findings highlighted the ability of large pretrained transformers to capture meaningful signals without project-specific tuning.
@@
Despite strong results, state-of-the-art text-based methods often require large amounts of labeled, project-specific data, which challenges smaller projects (zhou2025issuecourier; arnob2025bug). They also depend on titles and descriptions whose style and availability vary across projects (akhavan2025linkanchorautonomousllmbasedagent; zhang2023ealink), limiting generalizability. By fine-tuning a pretrained LLM we reduce the need for project-specific data, and by using standardized instruction prompts we lessen sensitivity to cross-project writing variation.
@@ 7.2. Graph-Based Methods
To address the limits of text-only models, recent research has explored graph-based methods that model the relationships between issues, developers, and other relevant entities. These methods are based on the idea that a developer’s history of resolved issues reflects their expertise (zhou2025issuecourier; wu2022spatial; cao2025complex; dai2023graph; dong2024neighborhood). Graph-based models build structures such as bipartite graphs or heterogeneous graphs to represent connections between reports, fixers, and code components. Graph Neural Networks (GNNs) (wu2020comprehensive) are then used to learn node embeddings that capture both the structure and content of the data.
@@
Dai et al. (dai2023graph) proposed GCBT, which constructed a bipartite graph linking issues and developers. The model initialized issue nodes using a GRU-based text encoder and learned embeddings for both issues and developers through GNNs. It then applied an information retrieval mechanism to rank candidates. Dong et al. (dong2024neighborhood) extended this approach in NCGBT by incorporating contrastive learning, which enriched node representations by leveraging both semantic and structural neighborhood information. Tao et al. (tao2025structural) introduced SCL-BT, a self-supervised framework that uses edge perturbation and hypergraph sampling to capture both labeled and unlabeled patterns in issue–developer graphs.
@@
Some recent work also incorporated temporal aspects of developer behavior (zhou2025issuecourier; wu2022spatial; cao2025complex). As contributors shift roles or become inactive, capturing these changes over time becomes critical. Wu et al. (wu2022spatial) developed ST-DGNN, a model that used joint random walk sampling and recurrent graph convolution to learn patterns in developer activity at hourly, daily, and weekly scales. Zhou et al. (zhou2025issuecourier) introduced IssueCourier, a multi-relational temporal GNN that modeled five types of relationships among issues, developers, and code files, and sliced the graph over time to track evolving team dynamics. Cao et al. (cao2025complex) focused on modeling inter-issue dependencies, representing issues as nodes and blocking relationships as edges. Their method divided the graph into time-ordered snapshots and trained a ranking model incrementally to assign issues based on recent changes.
@@
Although graph-based approaches offer strong modeling capabilities, they also present practical challenges. Many early methods struggled with scalability and high computational demands, which led researchers to adopt random walk sampling rather than full-graph learning (wu2022spatial; grover2016node2vec; nguyen2018continuous). While this reduced complexity, it also risked missing important patterns. Additionally, these models rely on explicit links between issues and developers, which are often sparse or noisy. This reliance can degrade model performance, especially in projects with incomplete or inconsistent tracking data (tao2025structural; dong2024neighborhood).
@@ 8. Conclusions and Future Work
While the results are promising, several avenues for future research remain. First, our evaluation was limited to two datasets; extending the analysis to a broader range of projects and domains would strengthen claims about generalizability. Second, incorporating retrieval-augmented generation (RAG) (lewis2020retrieval) techniques could enable the model to dynamically access relevant historical data, such as past issue–assignee pairs or commit histories, potentially boosting performance, as RAG-based approaches have shown strong results in other software maintenance tasks (huang2025back). Third, this study focused on fine-tuning a single LLM; future work should explore additional models and architectures to better understand their strengths and weaknesses for issue assignment. Forth, incorporating richer contextual information, such as developer profiles, project metadata, or historical collaboration patterns, into prompts could further improve assignment accuracy. Finally, beyond accuracy metrics, future research should evaluate the downstream impact of fine-tuning on developer productivity, issue resolution times, and integration into real-world workflows.

### v2

@@ 5. Discussion
Another key finding is that LIA consistently outperformed state-of-the-art graph-based approaches, including NCGBT (Dong et al., 2024). This is notable because graph-based methods have traditionally been strong baselines for issue assignment due to their ability to leverage project structure and developer–issue relationships. In addition, LIA also significantly outperformed our text-based baseline (CBR), showing that a fine-tuned LLM can capture deeper relational patterns between developers and issue semantics, even without relying on explicit graph structures.
@@ 6. Threats to Validity
We provide the entire experiment, including the source code for LIA and scripts evaluation online (41).
@@ 7. Related Work
Automatic issue assignment has become an important component in modern software maintenance, aiming to reduce manual effort and speed up the resolution of reported defects (Dipongkor and Moran, 2023; Lee et al., 2022). Over time, researchers have proposed a variety of techniques to automate this process. In this section, we group previous work into two categories: text classification-based and graph-based methods.
@@ 7.1. Text-Based Methods
Text-based approaches have commonly been used for issue assignment by treating parts of issue reports, such as titles and descriptions, as structured input. These inputs are turned into feature representations, and various models are trained to predict the appropriate developer as a classification label. Early work in this space used traditional ML techniques such as Naive Bayes (Murphy and Cubranic, 2004) and Support Vector Machines (SVM) (Anvik, 2006), coupled with methods like TF-IDF (Xuan et al., 2014) and Latent Dirichlet Allocation (LDA) (Xia et al., 2016). These models, however, lacked the semantic understanding needed to capture the complex context often found in issue report text (Cao and Cui, 2025; Zhou et al., 2025; Zhang et al., 2023; Arnob et al., 2025).
@@
With the development of DL, various classification techniques have been introduced to enhance issue assignment. Wang et al. (Wang et al., 2024) compared 35 DL-based issue assignment models, combining different text embeddings with various classifiers. They reported that Bi-LSTM models using attention or ELMo embeddings consistently outperformed alternatives in both top-k accuracy and mean reciprocal rank (MRR). Similarly, Mani et al. (Mani et al., 2019) developed a model combining a bidirectional recurrent neural network with attention, using Word2Vec (Mikolov, 2013) to represent text and learn both syntactic and semantic features. Aung et al. (Aung et al., 2022) introduced a multi-task learning model that simultaneously assigned developers and classified issue types. Their model used a CNN for encoding issue text and a BiLSTM for abstract syntax trees (ASTs), allowing it to incorporate structural code features often ignored by previous approaches. Arnob et al. (Arnob et al., 2025) used XLNet for issue assignment and demonstrated that combining commit messages with summaries and descriptions led to better performance by providing richer context than text alone.
@@
More recent studies have applied transformer models that are better at capturing contextual and semantic patterns than the mentioned DL approaches. Lee et al. (Lee et al., 2022) proposed a lightweight framework that fine-tunes pretrained transformers like RoBERTa (Liu et al., 2019) and DeBERTa (He et al., 2020) on issue report data, achieving competitive performance with lower computational overhead. Dipongkor et al. (Dipongkor and Moran, 2023) evaluated transformer-based models across multiple open-source datasets and found that DeBERTa outperformed BERT (Devlin et al., 2019) and RoBERTa due to its deeper attention mechanisms and disentangled representations. Their findings highlighted the ability of large pretrained transformers to capture meaningful signals without project-specific tuning.
@@
Despite strong results, state-of-the-art text-based methods often require large amounts of labeled, project-specific data, which challenges smaller projects (Zhou et al., 2025; Arnob et al., 2025). They also depend on titles and descriptions whose style and availability vary across projects (Akhavan et al., 2025; Zhang et al., 2023), limiting generalizability. By fine-tuning a pretrained LLM we reduce the need for project-specific data, and by using standardized instruction prompts we lessen sensitivity to cross-project writing variation.
@@ 7.2. Graph-Based Methods
To address the limits of text-only models, recent research has explored graph-based methods that model the relationships between issues, developers, and other relevant entities. These methods are based on the idea that a developer’s history of resolved issues reflects their expertise (Zhou et al., 2025; Wu et al., 2022; Cao and Cui, 2025; Dai et al., 2023; Dong et al., 2024). Graph-based models build structures such as bipartite graphs or heterogeneous graphs to represent connections between reports, fixers, and code components. Graph Neural Networks (GNNs) (Wu et al., 2020) are then used to learn node embeddings that capture both the structure and content of the data.
@@
Dai et al. (Dai et al., 2023) proposed GCBT, which constructed a bipartite graph linking issues and developers. The model initialized issue nodes using a GRU-based text encoder and learned embeddings for both issues and developers through GNNs. It then applied an information retrieval mechanism to rank candidates. Dong et al. (Dong et al., 2024) extended this approach in NCGBT by incorporating contrastive learning, which enriched node representations by leveraging both semantic and structural neighborhood information. Tao et al. (Tao et al., 2025) introduced SCL-BT, a self-supervised framework that uses edge perturbation and hypergraph sampling to capture both labeled and unlabeled patterns in issue–developer graphs.
@@
Some recent work also incorporated temporal aspects of developer behavior (Zhou et al., 2025; Wu et al., 2022; Cao and Cui, 2025). As contributors shift roles or become inactive, capturing these changes over time becomes critical. Wu et al. (Wu et al., 2022) developed ST-DGNN, a model that used joint random walk sampling and recurrent graph convolution to learn patterns in developer activity at hourly, daily, and weekly scales. Zhou et al. (Zhou et al., 2025) introduced IssueCourier, a multi-relational temporal GNN that modeled five types of relationships among issues, developers, and code files, and sliced the graph over time to track evolving team dynamics. Cao et al. (Cao and Cui, 2025) focused on modeling inter-issue dependencies, representing issues as nodes and blocking relationships as edges. Their method divided the graph into time-ordered snapshots and trained a ranking model incrementally to assign issues based on recent changes.
@@
Although graph-based approaches offer strong modeling capabilities, they also present practical challenges. Many early methods struggled with scalability and high computational demands, which led researchers to adopt random walk sampling rather than full-graph learning (Wu et al., 2022; Grover and Leskovec, 2016; Nguyen et al., 2018). While this reduced complexity, it also risked missing important patterns. Additionally, these models rely on explicit links between issues and developers, which are often sparse or noisy. This reliance can degrade model performance, especially in projects with incomplete or inconsistent tracking data (Tao et al., 2025; Dong et al., 2024).
@@ 8. Conclusions and Future Work
While the results are promising, several avenues for future research remain. First, our evaluation was limited to two datasets; extending the analysis to a broader range of projects and domains would strengthen claims about generalizability. Second, incorporating retrieval-augmented generation (RAG) (Lewis et al., 2020) techniques could enable the model to dynamically access relevant historical data, such as past issue–assignee pairs or commit histories, potentially boosting performance, as RAG-based approaches have shown strong results in other software maintenance tasks (Huang et al., 2025a). Third, this study focused on fine-tuning a single LLM; future work should explore additional models and architectures to better understand their strengths and weaknesses for issue assignment. Forth, incorporating richer contextual information, such as developer profiles, project metadata, or historical collaboration patterns, into prompts could further improve assignment accuracy. Finally, beyond accuracy metrics, future research should evaluate the downstream impact of fine-tuning on developer productivity, issue resolution times, and integration into real-world workflows.